Session: Nov 1998 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Tue Nov 17 09:38:08 MET 1998) Success
No known problems

Onsala (A. Polatidis at Tue Nov 17 11:19:24 MET 1998) Success
No known problems

Torun (K.Borkowski at Tue Nov 17 13:22:09 MET 1998) Minor failures
The LCP receiver continued to be unstable. Otherwise observations successful.

Effelsberg (dgraham at Tue Nov 17 16:07:01 MET 1998) Minor failures
due to multiple schedule versions after deadline and communication problems with PI, we observed partly with a schedule version which only recorded LCP.

Jodrell (Alastair Gunn at Wed Nov 18 11:30:42 MET 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Tue Nov 24 19:09:10 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: es026.antab.Read the comments in the file to get further information.